# Normal approximations for the multivariate inverse Gaussian distribution and asymmetric kernel smoothing on $d$-dimensional half-spaces

### v4

@@ 1.1 Overview of the literature
A more advantageous form of multivariate inverse Gaussian (MIG) distribution was proposed by Minami (2003), which is defined on $d$-dimensional half-spaces and is based on a reduced inverse relationship with multivariate Gaussian distributions; see Definition 1.1 below for its density function. It allows for easy derivations of its mean vector, covariance matrix, and higher-order cumulants. It is also reproductive, infinitely divisible, closed under linear transformations, and occurs as a limiting case of the multivariate Lagrange distribution (Minami, 2007). Furthermore, given its representation as the location of a $d$-variate Brownian motion with correlated components upon hitting a specified hyperplane for the first time, it is particularly straightforward to simulate.
@@ 1.2 Density estimation and boundary bias on half-spaces
The boundary bias problem refers to the systematic underestimation of density values near the boundary of the support, caused by the spill-over effect of fixed, symmetric kernels that fail to adapt their shape near boundaries. When the estimation point lies close to the boundary, a substantial portion of the kernel’s mass extends outside the support (into regions where the true density is zero), reducing the effective weight assigned to points within the support and leading to biased estimates. For example, the commonly used Gaussian kernel, (e.g., Chacón and Duong, 2018, p. 15), spreads its mass around the estimation point without accounting for the support’s constraints. This mismatch causes a portion of the kernel’s contribution to vanish outside the valid region, and the closer the estimation point is to the boundary, the larger the proportion of the kernel’s mass that lies outside the support, further amplifying the bias.
@@ 2.1 The univariate case
Consider the most common parametrization of the inverse Gaussian distribution, as defined, e.g., in Chapter 15 of Johnson, Kotz and Balakrishnan (1994). For any positive reals $\mu,\lambda\in(0,\infty)$, a random variable $X$ is said to be $\mathrm{IG}(\mu,\lambda)$-distributed, and one writes $X\sim\mathrm{IG}(\mu,\lambda)$, if its density function is given, for all $x\in(0,\infty)$, by
@@
An alternative elementary method based on Taylor expansions and Stirling’s formula was introduced by Cressie (1978) to obtain local approximations for the binomial distribution, along with refined continuity corrections. Non-asymptotic extensions of those results were provided by Ouimet (2022b) to sharpen Tusnády’s inequality in the “bulk” regime, an important ingredient in the Komlós–Major–Tusnády approximation controlling deviations between the empirical process and a corresponding sequence of Brownian bridges. Cressie’s line of reasoning was also adapted by Ouimet (2023) to study the asymptotic behavior of the median of the negative binomial distribution jittered by a uniform, settling an open question of Cœurjolly and Trépanier (2020). The proof of Theorem 2.1 employs a similar elementary approach; see Appendix A.1 for details.
@@ 2.2 The multivariate case
where $\smash{\mathbf{\Omega}_{0}^{1/2}}$ denotes the symmetric square root of $\mathbf{\Omega}_{0}$. This is a straightforward consequence of the expression for the MIG cumulant generating function in Theorem 1 of Minami (2003) combined with Lévy’s continuity theorem for Laplace transforms; see, e.g., (Feller, 1971, p. 431).
@@
In particular, as long as $\boldsymbol{x}$ is located in the bulk of the MIG distribution (i.e., not too far from the mean $\mu\boldsymbol{\xi}_{0}$), the terms in the series expansion of the log-ratio $\mathrm{LR}(\boldsymbol{x})$ will be shown explicitly in Theorem 2.3, along with the rate of convergence of any truncation. This result rests indirectly on the fact that the MIG distribution is infinitely divisible, as stated in Property 4 of Minami (2003).
@@
(a) $\mathbf{\Omega}_{0}=\begin{pmatrix}2&1\\ 1&2\end{pmatrix}$
(b) $\mathbf{\Omega}_{0}=\begin{pmatrix}2&1\\ 1&3\end{pmatrix}$
(c) $\mathbf{\Omega}_{0}=\begin{pmatrix}2&1\\ 1&4\end{pmatrix}$
(d) $\mathbf{\Omega}_{0}=\begin{pmatrix}3&1\\ 1&2\end{pmatrix}$
(e) $\mathbf{\Omega}_{0}=\begin{pmatrix}3&1\\ 1&3\end{pmatrix}$
(f) $\mathbf{\Omega}_{0}=\begin{pmatrix}3&1\\ 1&4\end{pmatrix}$
(g) $\mathbf{\Omega}_{0}=\begin{pmatrix}4&1\\ 1&2\end{pmatrix}$
(h) $\mathbf{\Omega}_{0}=\begin{pmatrix}4&1\\ 1&3\end{pmatrix}$
(i) $\mathbf{\Omega}_{0}=\begin{pmatrix}4&1\\ 1&4\end{pmatrix}$
@@
(a) $\mathbf{\Omega}_{0}=\begin{pmatrix}2&1\\ 1&2\end{pmatrix}$
(b) $\mathbf{\Omega}_{0}=\begin{pmatrix}2&1\\ 1&3\end{pmatrix}$
(c) $\mathbf{\Omega}_{0}=\begin{pmatrix}2&1\\ 1&4\end{pmatrix}$
(d) $\mathbf{\Omega}_{0}=\begin{pmatrix}3&1\\ 1&2\end{pmatrix}$
(e) $\mathbf{\Omega}_{0}=\begin{pmatrix}3&1\\ 1&3\end{pmatrix}$
(f) $\mathbf{\Omega}_{0}=\begin{pmatrix}3&1\\ 1&4\end{pmatrix}$
(g) $\mathbf{\Omega}_{0}=\begin{pmatrix}4&1\\ 1&2\end{pmatrix}$
(h) $\mathbf{\Omega}_{0}=\begin{pmatrix}4&1\\ 1&3\end{pmatrix}$
(i) $\mathbf{\Omega}_{0}=\begin{pmatrix}4&1\\ 1&4\end{pmatrix}$
@@ 3.1 Asymptotic results
In this section, the local limit theorem for the MIG distribution (Theorem 2.3) is applied to derive the asymptotic variance of the new density estimator $\smash{\hat{f}_{n,\mathbf{H}}}$ at each point $\boldsymbol{\xi}\in\mathcal{H}(\boldsymbol{\beta})$; see Proposition 3.2. Combined with a careful analysis of the asymptotic bias (Proposition 3.1), these results yield the leading-order expansions for both the mean squared error (MSE) (Corollary 3.3) and the MISE (Corollary 3.4) and enable the derivation of asymptotically optimal bandwidths. Moreover, the asymptotic normality of $\smash{\hat{f}_{n,\mathbf{H}}}$ at each point $\boldsymbol{\xi}$ is established in Theorem 3.5. The proofs are deferred to Appendix B.
@@
$\mathsf{Bias}\big{\{}\hat{f}_{n,\mathbf{H}}(\boldsymbol{\xi})\big{\}}$ $=\frac{1}{2}\left\{\sum_{\ell=1}^{d}\beta_{\ell}c_{\ell}\lambda_{\ell}\right\}\sum_{k=1}^{d}\lambda_{k}\left\{\sum_{i,j=1}^{d}\frac{v_{ik}v_{jk}}{c_{i}c_{j}}\frac{\partial^{2}f(\boldsymbol{\xi})}{\partial\lambda_{i}\partial\lambda_{j}}\right\}$
$=\mathcal{O}(\|\mathbf{H}\|_{2}^{2}).$
@@
Although $\smash{\hat{f}_{n,\mathbf{H}}}$ exhibits superior boundary bias behavior, this benefit comes at the cost of increasing the variance near the boundary, as the next proposition reveals. This tradeoff is typical for asymmetric kernel estimators; see, e.g., Chen (1999) or Ouimet and Tolosana-Delgado (2022). It is also a recurring feature in the study of Bernstein estimators; see, e.g., Leblanc (2012) or Ouimet (2022f).
@@
In particular, if all eigenvalues of $\mathbf{H}$ stay equal as $n\to\infty$, then the variance of the estimator near the origin is $\mathcal{O}(n^{-1}|\mathbf{H}|^{-1})$.
@@
The asymptotic rate of the optimal MISE, $n^{-4/(d+4)}$, established in Corollary 3.4, is known to be minimax for second-order kernel density estimators under the assumption that the second-order partial derivatives of the target density are uniformly continuous and bounded; see, e.g., Farrell (1972), Stone (1980) or Tsybakov (2009). For a detailed treatment of the minimaxity of the related Dirichlet kernel density estimator on the simplex, considering a wide range of $L^{p}$ loss functions and degrees of smoothness of the target density, refer to Bertin et al. (2023). Extending their analysis to the present setting is an interesting direction for future research.
@@ 3.2.1 Competing estimators
where $\Phi(\,\cdot\,;\mu,\sigma^{2})$ denotes the univariate cdf of the $\mathcal{N}(\mu,\sigma^{2})$ distribution. Note that, since $\boldsymbol{\xi}\in\mathcal{H}(\boldsymbol{\beta})$, the mode of every truncated Gaussian component in the sum necessarily lies within the half-space.
@@ 3.2.3 Bandwidth selection
$\hat{f}^{-(i)}_{n-1,\mathbf{H}}(\boldsymbol{x}_{i})=\frac{1}{n-1}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}k_{\boldsymbol{\beta},\boldsymbol{x}_{i},\mathbf{H}}(\boldsymbol{x}_{j}).$
See, e.g., Section 2.1 of Zhang, King and Hyndman (2006) for more details. The latter method is sensitive to extremes and outliers and may lead to oversmoothing when such observations are present. Other options include the least-squares cross-validation (LSCV) criterion (Rudemo, 1982; Bowman, 1984), viz.
@@
using the automatic threshold proposed by Wu (2019), namely
@@
with
$\widehat{\boldsymbol{\Sigma}}_{n}=\frac{1}{n}\sum_{i=1}^{n}(\boldsymbol{x}_{i}-\bar{\boldsymbol{x}}_{n})(\boldsymbol{x}_{i}-\bar{\boldsymbol{x}}_{n})^{\top}$
denoting the empirical covariance matrix.
@@
Estimators that minimize the AMISE (MIG kernel only) as well as those optimizing the LCV, LSCV, and RLCV criteria were examined. With the exception of the transformation estimator, $\smash{\widehat{f}_{n,\mathbf{H}}^{(\mathrm{trans})}(\boldsymbol{\xi})}$, the entries of the unstructured $d\times d$ bandwidth matrix were optimized using a parametrization based on its Cholesky root with positivity constraints on the diagonal (Pinheiro and Bates, 1996). Additionally, the isotropic AMISE bandwidth matrix of the form $h^{2}\mathbf{S}_{n}$ based on the transformed data, as discussed in Remark 3.5, was considered.
@@ 3.2.4 Performance measures
Both the RMISE and BRMISE are estimated via numerical integration using the $h$-adaptive integration routines of Steven G. Johnson (Genz and Malik, 1980; Berntsen, Espelid and Genz, 1991; Johnson, 2005) after rotating the space with the projection matrix $\mathbf{Q}$ from Proposition 1.2, with an absolute error tolerance of $10^{-4}$ or a maximum of $5\times 10^{5}$ evaluations.
@@ 3.2.5 Results
    No combinations of kernel and bandwidth selection method universally dominate the others for any of the performance metrics considered.
@@ 3.2.6 Discussion
    Complexity: The LCV criterion is $\mathcal{O}(n)$ and the number of bandwidth entries to optimize for a $d$-dimensional model is $\mathcal{O}(d^{2})$, with $d(d+1)/2$ constraints. As expected, the spectral norm of the optimal bandwidth matrix increases with the dimension $d$ and decreases as the sample size $n$ grows. Using a diagonal bandwidth matrix reduces to $d$ the number of parameters to be estimated, and leads to gains in speed at the expense of poorer performance (higher MISE).
@@ 3.3 Real-data application
Coupled with a maximum data information prior, $10^{3}$ draws from the posterior distribution of $(\sigma_{u},\xi)$ are obtained, supported on the (truncated) half-space
@@
The left panel of Figure 9 shows the $10^{3}$ posterior draws used for estimating the kernel density. Given the small sample size, $n=52$, the posterior distribution of the generalized Pareto is far from Gaussian, notwithstanding the support constraint. The graph includes a scatterplot of the observations and contour curves for the MIG density evaluated at the maximum likelihood estimate.

### v5

@@ 1.1 Overview of the literature
A more advantageous form of multivariate inverse Gaussian (MIG) distribution was proposed by Minami (2003). The latter is defined on $d$-dimensional half-spaces and is based on a reduced inverse relationship with multivariate Gaussian distributions; see Definition 1.1 below for its density. It allows for easy derivations of its mean vector, covariance matrix, and higher-order cumulants. It is also reproductive, infinitely divisible, closed under linear transformations, and occurs as a limiting case of the multivariate Lagrange distribution (Minami, 2007). Furthermore, given its representation as the location of a $d$-variate Brownian motion with correlated components upon hitting a specified hyperplane for the first time, it is straightforward to simulate.
@@ 1.2 Density estimation and boundary bias on half-spaces
The boundary bias problem refers to the systematic underestimation of density values near the boundary of the support, caused by the spill-over effect of fixed, symmetric kernels that fail to adapt their shape near boundaries. When the estimation point lies close to the boundary, a substantial portion of the kernel’s mass extends outside the support (into regions where the true density is zero), reducing the effective weight assigned to points within the support and leading to biased estimates. For example, the commonly used Gaussian kernel (e.g., Chacón and Duong, 2018, p. 15) spreads its mass around the estimation point without accounting for the support’s constraints. This mismatch causes a portion of the kernel’s contribution to vanish outside the valid region, and the closer the estimation point is to the boundary, the larger the proportion of the kernel’s mass that lies outside the support, further amplifying the bias.
@@ 2.1 The univariate case
Consider the most common parametrization of the inverse Gaussian distribution, as defined, e.g., in Chapter 15 of Johnson, Kotz and Balakrishnan (1994). For any positive reals $\mu,\lambda\in(0,\infty)$, a random variable $X$ is said to be $\mathrm{IG}(\mu,\lambda)$-distributed, and one writes $X\sim\mathrm{IG}(\mu,\lambda)$, if its density is given, for all $x\in(0,\infty)$, by
@@
An alternative elementary method based on Taylor expansions and Stirling’s formula was introduced by Cressie (1978) to obtain local approximations for the binomial distribution, along with refined continuity corrections. Non-asymptotic extensions of those results were provided by Ouimet (2022b) to sharpen Tusnády’s inequality in the “bulk” regime, an important ingredient in the Komlós–Major–Tusnády approximation controlling deviations between the empirical process and a corresponding sequence of Brownian bridges. Cressie’s line of reasoning was also adapted by Ouimet (2023) to study the asymptotic behavior of the median of the negative binomial distribution jittered by a uniform, settling an open question of Cœurjolly and Trépanier (2020). The proof of Theorem 2.1 uses a similar elementary approach; see Appendix A.1 for details.
@@ 2.2 The multivariate case
where $\smash{\mathbf{\Omega}_{0}^{1/2}}$ denotes the symmetric square root of $\mathbf{\Omega}_{0}$. This is a straightforward consequence of the expression for the MIG cumulant generating function in Theorem 1 of Minami (2003) combined with Lévy’s continuity theorem for Laplace transforms; see, e.g., Feller (1971, p. 431).
@@
In particular, as long as $\boldsymbol{x}$ is located in the bulk of the MIG distribution (i.e., not too far from the mean $\mu\boldsymbol{\xi}_{0}$), the terms in the series expansion of the log-ratio $\mathrm{LR}(\boldsymbol{x})$ will be shown explicitly in Theorem 2.3, along with the rate of convergence of any truncation. This result rests indirectly on the fact that the MIG distribution is infinitely divisible, as per Property 4 of Minami (2003).
@@
(a) $\mathbf{\Omega}_{0}=\left(\begin{smallmatrix}2&1\\ 1&2\end{smallmatrix}\right)$
(b) $\mathbf{\Omega}_{0}=\left(\begin{smallmatrix}2&1\\ 1&3\end{smallmatrix}\right)$
(c) $\mathbf{\Omega}_{0}=\left(\begin{smallmatrix}2&1\\ 1&4\end{smallmatrix}\right)$
(d) $\mathbf{\Omega}_{0}=\left(\begin{smallmatrix}3&1\\ 1&2\end{smallmatrix}\right)$
(e) $\mathbf{\Omega}_{0}=\left(\begin{smallmatrix}3&1\\ 1&3\end{smallmatrix}\right)$
(f) $\mathbf{\Omega}_{0}=\left(\begin{smallmatrix}3&1\\ 1&4\end{smallmatrix}\right)$
(g) $\mathbf{\Omega}_{0}=\left(\begin{smallmatrix}4&1\\ 1&2\end{smallmatrix}\right)$
(h) $\mathbf{\Omega}_{0}=\left(\begin{smallmatrix}4&1\\ 1&3\end{smallmatrix}\right)$
(i) $\mathbf{\Omega}_{0}=\left(\begin{smallmatrix}4&1\\ 1&4\end{smallmatrix}\right)$
@@
(a) $\mathbf{\Omega}_{0}=\left(\begin{smallmatrix}2&1\\ 1&2\end{smallmatrix}\right)$
(b) $\mathbf{\Omega}_{0}=\left(\begin{smallmatrix}2&1\\ 1&3\end{smallmatrix}\right)$
(c) $\mathbf{\Omega}_{0}=\left(\begin{smallmatrix}2&1\\ 1&4\end{smallmatrix}\right)$
(d) $\mathbf{\Omega}_{0}=\left(\begin{smallmatrix}3&1\\ 1&2\end{smallmatrix}\right)$
(e) $\mathbf{\Omega}_{0}=\left(\begin{smallmatrix}3&1\\ 1&3\end{smallmatrix}\right)$
(f) $\mathbf{\Omega}_{0}=\left(\begin{smallmatrix}3&1\\ 1&4\end{smallmatrix}\right)$
(g) $\mathbf{\Omega}_{0}=\left(\begin{smallmatrix}4&1\\ 1&2\end{smallmatrix}\right)$
(h) $\mathbf{\Omega}_{0}=\left(\begin{smallmatrix}4&1\\ 1&3\end{smallmatrix}\right)$
(i) $\mathbf{\Omega}_{0}=\left(\begin{smallmatrix}4&1\\ 1&4\end{smallmatrix}\right)$
@@ 3.1 Asymptotic results
In this section, the local limit theorem for the MIG distribution (Theorem 2.3) is applied to derive the asymptotic variance of the new density estimator $\smash{\hat{f}_{n,\mathbf{H}}}$ at each point $\boldsymbol{\xi}\in\mathcal{H}(\boldsymbol{\beta})$; see Proposition 3.2. Combined with a careful analysis of the asymptotic bias (Proposition 3.1), this result yields the leading-order expansions for both the mean squared error (MSE) (Corollary 3.3) and the MISE (Corollary 3.4) and enable the derivation of asymptotically optimal bandwidths. Moreover, the asymptotic normality of $\smash{\hat{f}_{n,\mathbf{H}}}$ at each point $\boldsymbol{\xi}$ is established in Theorem 3.5. The proofs are deferred to Appendix B.
@@
$\mathsf{Bias}\big{\{}\hat{f}_{n,\mathbf{H}}(\boldsymbol{\xi})\big{\}}=\frac{1}{2}\left\{\sum_{\ell=1}^{d}\beta_{\ell}c_{\ell}\lambda_{\ell}\right\}\sum_{k=1}^{d}\lambda_{k}\left\{\sum_{i,j=1}^{d}\frac{v_{ik}v_{jk}}{c_{i}c_{j}}\frac{\partial^{2}f(\boldsymbol{\xi})}{\partial\lambda_{i}\partial\lambda_{j}}\right\}=\mathcal{O}(\|\mathbf{H}\|_{2}^{2}).$
@@
Although $\smash{\hat{f}_{n,\mathbf{H}}}$ exhibits superior boundary bias behavior, this benefit comes at the cost of increasing the variance near the boundary, as the next proposition reveals. This phenomenon is a direct consequence of the adaptive mechanism of the asymmetric MIG kernel. To avoid the boundary bias problem that affects symmetric kernels, the MIG kernel adapts its shape based on the estimation point $\xi$. As $\xi$ approaches the boundary of the half-space (i.e., as $\boldsymbol{\beta}^{\top}\boldsymbol{\xi}\to 0$), the kernel becomes increasingly skewed and concentrated to confine its mass within the support. This necessary adaptation ensures that the bias remains low. However, by becoming more “peaked,” the resulting density estimate at that point relies more heavily on a smaller effective number of local data points. This heightened sensitivity to the specific locations of a few nearby observations leads to greater sampling variability, and thus, a higher variance. This behavior is typical for asymmetric kernel estimators; see, e.g., Chen (1999) or Ouimet and Tolosana-Delgado (2022). It is also a recurring feature in the study of Bernstein estimators; see, e.g., Leblanc (2012) or Ouimet (2022f).
In contrast, a classical kernel uses a fixed, symmetric shape, which leads to some of the kernel’s mass to spill over and higher boundary bias. The variance of this type of estimator, however, is not structurally affected by the boundary. Given that the variance depends on the kernel’s fixed shape and the target density $f$, rather than the point’s proximity to the boundary, it remains stable across the domain. This illustrates the alternative tradeoff: putting up with high bias in exchange for stable variance.
@@
In particular, if all eigenvalues of $\mathbf{H}$ remain equal as $n\to\infty$, then the variance of the estimator near the origin is $\mathcal{O}(n^{-1}|\mathbf{H}|^{-1})$.
@@
The asymptotic rate of the optimal MISE, $n^{-4/(d+4)}$, established in Corollary 3.4, is known to be minimax for second-order kernel density estimators under the assumption that the second-order partial derivatives of the target density are uniformly continuous and bounded; see, e.g., Farrell (1972), Stone (1980) or Tsybakov (2009). For a detailed treatment of the minimaxity of the related Dirichlet kernel density estimator on the simplex, considering a wide range of $L^{p}$ loss functions and degrees of smoothness of the target density, refer to Bertin et al. (2023). Extending these authors’ analysis to the present setting is an interesting direction for future research.
@@ 3.2.1 Competing estimators
where $\Phi(\,\cdot\,;\mu,\sigma^{2})$ denotes the univariate cdf of the $\mathcal{N}(\mu,\sigma^{2})$ distribution. Note that, because $\boldsymbol{\xi}\in\mathcal{H}(\boldsymbol{\beta})$, the mode of every truncated Gaussian component in the sum necessarily lies within the half-space.
@@ 3.2.3 Bandwidth selection
$\hat{f}^{-(i)}_{n-1,\mathbf{H}}(\boldsymbol{x}_{i})=\frac{1}{n-1}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}k_{\boldsymbol{\beta},\boldsymbol{x}_{i},\mathbf{H}}(\boldsymbol{x}_{j});$
see, e.g., Section 2.1 of Zhang, King and Hyndman (2006) for more details. The LCV criterion is sensitive to extremes and outliers and may lead to oversmoothing when such observations are present. Other options include the least-squares cross-validation (LSCV) criterion (Rudemo, 1982; Bowman, 1984), viz.
@@
using the automatic threshold proposed by Wu (2019),
@@
with $\widehat{\boldsymbol{\Sigma}}_{n}=n^{-1}\sum_{i=1}^{n}(\boldsymbol{x}_{i}-\bar{\boldsymbol{x}}_{n})(\boldsymbol{x}_{i}-\bar{\boldsymbol{x}}_{n})^{\top}$ denoting the empirical covariance matrix.
@@
Estimators that minimize the AMISE (MIG kernel only) as well as those optimizing the LCV, LSCV, and RLCV criteria were examined. With the exception of the transformation estimator, $\smash{\widehat{f}_{n,\mathbf{H}}^{(\mathrm{trans})}(\boldsymbol{\xi})}$, the entries of the unstructured $d\times d$ bandwidth matrix were optimized using a parametrization based on its Cholesky root with positivity constraints on the diagonal (Pinheiro and Bates, 1996). Also considered was the isotropic AMISE bandwidth matrix of the form $h^{2}\mathbf{S}_{n}$ based on the transformed data, as discussed in Remark 3.5.
@@ 3.2.4 Performance measures
Both the RMISE and BRMISE are estimated via numerical integration using the $h$-adaptive integration routines of Steven G. Johnson (Genz and Malik, 1980; Berntsen, Espelid and Genz, 1991; Johnson, 2005) after rotating the space with the projection matrix $\mathbf{Q}$ from Proposition 1.2, with an absolute error tolerance of $10^{-4}$ or up to a maximum of $5\times 10^{5}$ evaluations.
@@ 3.2.5 Results
    No combination of kernel and bandwidth selection method universally dominates the others for any of the performance metrics considered.
@@ 3.2.6 Discussion
    Complexity: The LCV criterion is $\mathcal{O}(n)$ and the number of bandwidth entries to optimize for a $d$-dimensional model is $d(d+1)/2$, so $\mathcal{O}(d^{2})$. As expected, the spectral norm of the optimal bandwidth matrix increases with the dimension $d$ and decreases as the sample size $n$ grows. Using a diagonal bandwidth matrix reduces to $d$ the number of parameters to be estimated, and leads to gains in speed at the expense of poorer performance (higher MISE).
@@ 3.3 Real-data application
Coupled with a maximum data information prior, $10^{3}$ independent draws from the posterior distribution of $(\sigma_{u},\xi)$ were obtained using the generalized ratio-of-uniform method (Northrop, 2024; Wakefield, Gelfand and Smith, 1991). The bivariate posterior is supported on the (truncated) half-space
@@
The left panel of Figure 9 shows the $10^{3}$ posterior draws used for estimating the kernel density. Given the small sample size, $n=52$, the posterior distribution of the generalized Pareto is far from Gaussian, notwithstanding the support constraint. The panel shows both a scatterplot of the observations and contour curves for the MIG density with maximum-likelihood-estimated parameters.